OpenStudy (anonymous):

Ahooy! Baby Anna weighed 96 ounces at birth and x ounces by the end of the first year. Baby Allison weighed 78 ounces at birth and gained twice as much weight as Baby Anna by the end of her first year. If at the end of the year both babies weighed the same amount, how much weight did Baby Allison gain?

OpenStudy (anonymous):

At the end of the year baby anna weighed 96+x oz. and baby allison weighed 78+2x oz. these two expressions are equal to each other. Solve for x. 96+x=78+2x -x -x 96=78+x -78 -78 18=x Baby Anna gained 18 oz, and Baby Allison gained 36 oz

OpenStudy (anonymous):

Great! thank you,